# Visual Self-Fulfilling Alignment: Shaping Safety-Oriented Personas via Threat-Related Images

### 09 Mar 2026 v1

@@ 4.3 Ablation Studies
Figure 2 presents the ablation on fine-tuning modality. We compare three variants: Image, Text, and Mixed. Image achieves the best results across all models. Why? Text training only modifies the LLM backbone. It leaves the visual pathway unchanged. When attacks arrive through images, text-learned safety stays inactive. Safety alignment in one modality does not transfer to another. Mixed training has a different problem. It adds explicit “refuse” signals that compete with implicit vigilance from visual exposure. This dilutes the self-fulfilling effect. Image works through a different mechanism. Threat-related content activates persona features tied to vigilance. The model internalizes caution through what it sees, not what it is told. Image beats Mixed by 4-7% ASR across all models. Even Text achieves higher CS than AdaShield. Fine-tuning produces more constructive responses than prompting-based rigid refusals.
@@ 5 Conclusion
We introduce VSFA, a method that extends the self-fulfilling mechanism from text to visual modalities. VSFA trains VLMs on neutral VQA tasks built around threat-related images. The training data contains no safety labels or contrastive pairs. Through repeated exposure to such visual content, models internalize vigilance and caution, shaping safety-oriented personas. Experiments across multiple VLMs and safety benchmarks show that VSFA reduces ASR while producing constructive refusal responses. The method preserves general capabilities with minimal over-refusal. Compared to prompting-based defenses, VSFA avoids rigid rejections. Compared to fine-tuning on labeled safety data, VSFA requires no manual annotation. Our findings suggest that self-fulfilling mechanism operates effectively in the visual modality, offering a label-free approach to VLMs alignment.
@@
This work assumes that visual exposure alone can shape model behavior. Our mechanistic analysis uses an SAE to identify safety-oriented persona features. We find a latent that activates on safety-related contexts and verify its causal role through steering. However, SAE-based interpretability has known limitations. Feature isolation may be incomplete. The training images are all synthetic. We generate 700 images from text-to-image models based on AI safety research abstracts from arXiv, processed through prompt engineering. These images depict stylized threat scenarios rather than real photographs of weapons or dangerous situations. The visual style also reflects specific cultural conventions. We test on four VLMs at the 7B-8B scale from two model families. We do not evaluate larger models. The three safety benchmarks focus on jailbreak attacks. Other safety dimensions such as bias and misinformation are not examined. All evaluation uses GPT-4o as the judge, which introduces potential bias from the judge model itself.

### 15 Apr 2026 v2

@@ 4.3 Ablation Studies
Figure 2 presents the ablation on the fine-tuning modality. We compare three variants: Image, Text, and Mixed. Image achieves the best results across all models. This is because text training only modifies the LLM backbone. It leaves the visual pathway unchanged. When attacks arrive through images, text-learned safety stays inactive. Safety alignment in one modality does not transfer to another. Mixed training has a different problem. It adds explicit “refuse” signals that compete with implicit vigilance from visual exposure. This dilutes the self-fulfilling effect. Image works through a different mechanism. Threat-related content activates persona features tied to vigilance. The model internalizes caution through what it sees, not what it is told. Image beats Mixed by 4-7% ASR across all models. Even Text achieves higher CS than AdaShield. Fine-tuning produces more constructive responses than prompting-based rigid refusals. We also conducted a visual style ablation across 12 styles, showing that the safety effect is style-invariant (Appendix E).
@@ 5 Conclusion
We introduce VSFA, a method that extends the self-fulfilling mechanism from text to visual modalities. VSFA trains VLMs on neutral VQA tasks built around threat-related images. The training data contains no safety labels or contrastive pairs. Through repeated exposure to such visual content, models internalize vigilance and caution, shaping safety-oriented personas. Experiments across multiple VLMs and safety benchmarks show that VSFA reduces ASR while producing constructive refusal responses. The method preserves general capabilities with minimal over-refusal. Compared to prompting-based defenses, VSFA avoids rigid rejections. Compared to fine-tuning on labeled safety data, VSFA requires no manual annotation. Our findings suggest that a self-fulfilling mechanism operates effectively in the visual modality, offering a label-free approach to VLMs alignment.
@@
This work assumes that visual exposure alone can shape model behavior. Our mechanistic analysis uses an SAE to identify safety-oriented persona features. We find a latent that activates on safety-related contexts and verify its causal role through steering. However, SAE-based interpretability has known limitations. Feature isolation may be incomplete. The training images are all synthetic. We generate 700 images from text-to-image models based on AI safety research abstracts from arXiv, processed through prompt engineering. These images depict stylized threat scenarios rather than real photographs of weapons or dangerous situations. The visual style also reflects specific cultural conventions. We test on four VLMs at the 7B-8B scale from two model families. We do not evaluate larger models. The three safety benchmarks focus on jailbreak attacks. Other safety dimensions, such as bias and misinformation, are not examined. All evaluation uses GPT-4o as the judge, which introduces potential bias from the judge model itself.
Acknowledgments
Lijie Hu is supported by the funding BF0100 from Mohamed bin Zayed University of Artificial Intelligence (MBZUAI). Di Wang and Shu Yang are supported in part by the funding BAS/1/1689-01-01,RGC/3/7125-01-01, FCC/1/5940-20-05, FCC/1/5940-06-02, and King Abdullah University of Science and Technology (KAUST) – Center of Excellence for Generative AI, under award number 5940 and a gift from Google.
@@ C.7 Summary
C.8 Cross-Model Evidence
Our SAE analysis focuses on Qwen2.5-VL-7B. A natural question is whether persona features exist in other model families. Our main experiments already show that VSFA produces consistent ASR reductions across four models from the Qwen and LLaVA families (Table 1). This suggests a shared mechanism. Independent work by Arditi et al. Arditi and Chen (2025) provides direct evidence. They trained SAEs on Llama-3.1-8B-Instruct and Qwen2.5-7B-Instruct and applied model-diffing between the original and emergently misaligned versions. They found interpretable persona features in both models. These features correspond to undesirable traits like toxicity and manipulation. The features were consistent across multiple random seeds. For Llama, all 10 misalignment-relevant features appeared across 3 different fine-tuning runs. For Qwen, 9 out of 10 appeared in at least 2 runs. This confirms that persona features are not specific to one architecture. They are a general property of instruction-tuned language models. Our work shows the other side of this coin. If misaligned persona features exist across model families, safety-oriented persona features should too. VSFA activates these safety features through visual exposure, and the consistent results across our six tested models support this view.
Appendix D Teacher Model Comparison
This appendix reports our comparison of four teacher models for VSFA data construction: GPT-4o-mini OpenAI (2023), GPT-5, Claude 4.5 Sonnet, and Gemini-3-pro. The teacher model handles concept extraction, image prompt generation, and VQA answer generation in our pipeline. To select the best one, we ran the full VSFA pipeline with each model separately. Each model produced prompts for 60 images and about 360 QA pairs. We used Doubao as the text-to-image model for all conditions, fine-tuned Qwen2.5-VL-7B-Instruct with LoRA under identical hyperparameters, and evaluated on FigStep using GPT-4o as the judge. Table 9 shows the results.
Table 9: FigStep ASR across teacher models on Qwen2.5-VL-7B. All models use 60 images and about 360 QA pairs. Lower ASR means safer.
Teacher Model Images QA Pairs Avg. Length ASR (%) $\downarrow$
Baseline (no fine-tuning) 35.6
GPT-4o-mini (ours) 60 360 659 chars 12.0
Claude 4.5 Sonnet 60 360 1,322 chars 22.2
Gemini-3-pro 60 360 2,467 chars 28.8
GPT-5 60 360 769 chars 22.8
GPT-4o-mini achieves 12.0% ASR, a 66% reduction from the 35.6% baseline. The three stronger models also reduce ASR but by much less. Claude 4.5 Sonnet and GPT-5 reach about 22%, and Gemini-3-pro only drops to 28.8%. Answer length does not explain this gap. GPT-5 and GPT-4o-mini produce answers of similar length, but GPT-5’s ASR is nearly twice as high. We looked at the image prompts from each model and found the real difference there. GPT-4o-mini writes short, concrete prompts. They focus on specific visual elements like dark lighting, red warning signs, and surveillance cameras. Doubao receives clear instructions from these prompts and produces images with strong, consistent threat atmosphere. The concise answers from GPT-4o-mini also train the student model to respond briefly. This helps the model produce short refusals at test time. The stronger models produce noisier prompts. GPT-5 tends toward abstract, high-level descriptions. Claude 4.5 Sonnet tries to balance threatening and reassuring elements in the same prompt. Gemini-3-pro packs too many visual details into each prompt. In all three cases, Doubao receives unfocused instructions and produces images with weaker threat signals. This pattern is consistent with observations in knowledge distillation research. Gu et al. Gu et al. (2024) showed that the largest model is not always the best teacher, and smaller models can sometimes produce more focused training signals. Based on this comparison, we selected GPT-4o-mini as the teacher model for all experiments in the main paper. Among the four models we tested, it produces the clearest threat imagery and the most effective training data.
Appendix E Visual Style Ablation
This appendix examines whether the safety effect of VSFA depends on any particular visual style. Our main experiments use a mix of 12 visual styles when generating threat-related images. To isolate the contribution of each style, we trained Qwen2.5-VL-7B separately with 50 images of each style. We generated the corresponding VQA pairs through the same pipeline described in Section 3. We used LoRA rank 128 and only fine-tuned the language model. The visual encoder stayed frozen. We evaluated each model on FigStep with GPT-4o as the judge. Table 10 shows the results.
Table 10: Visual style ablation on Qwen2.5-VL-7B. Each row is a separate training run using 50 images of one style. Lower ASR means safer.
Visual Style Images FigStep ASR (%) $\downarrow$
No Defense (baseline) 0 35.6
Photorealistic 50 12.8
Digital art 50 13.1
Concept art 50 12.4
Technical illustration 50 13.7
Documentary style 50 11.9
Abstract representation 50 12.6
Cinematic 50 11.4
Artistic 50 11.2
Professional 50 12.1
Casual 50 13.1
Futuristic 50 12.8
Vintage 50 13.3
All 12 styles achieve ASR between 11.2% and 13.7%. The mean is 12.5% and the standard deviation is only 0.74%. Even the highest and lowest ASR only differ by 2.5 percentage points. Photorealistic and Abstract representation are the two most visually different styles in our set. One looks like a photograph. The other is stylized and non-literal. Yet their ASR results are almost the same , a gap of just 0.2 percentage points. This means the safety effect does not depend on visual style. The alignment signal comes from what the image depicts, not from how it looks. Our SAE analysis in Appendix C supports this. The most activated latent encodes abstract safety concepts rather than visual style features. Our finding also has a practical benefit. Practitioners do not need to optimize for any particular visual style when building VSFA training sets. It also explains why VSFA generalizes to attack categories not seen during training. The safety persona is activated by the semantic content of threat-related scenes, regardless of how they are rendered.
Appendix F General Capability Evaluation (MMLU and MMMU)
This appendix reports VSFA’s performance on two general reasoning benchmarks: MMLU Hendrycks et al. (2021) and MMMU Yue et al. (2024). MMLU covers 57 subjects across STEM, humanities, social sciences, and other domains. MMMU covers 30 subjects with college-level questions that require both visual and textual reasoning. We evaluated Qwen2.5-VL-7B before and after VSFA training using the standard protocol for both benchmarks. Table 11 shows the results.
Table 11: General capability evaluation on Qwen2.5-VL-7B before and after VSFA training.
Benchmark Baseline VSFA Diff
MMLU 68.52% 68.13% -0.39%
MMMU 50.56% 50.11% -0.45%
MMLU drops by 0.39% and MMMU drops by 0.45%. Both drops are well under 0.5%, which is within the normal variance of safety fine-tuning methods. Zong et al. Zong et al. (2024) reported MMLU changes between -0.08% and +1.79% across different VLGuard configurations. Our drops are at the lower end of this range. Combined with MM-Vet results in Table 2, VSFA preserves general capabilities across MMLU, MMMU, and MM-Vet.
Appendix G Cross-Family Generalization
This appendix tests whether VSFA generalizes beyond the 7B to 8B models used in the main paper. Our main experiments cover four models from two families: Qwen2.5-VL-7B, Qwen3-VL-8B, LLaVA-1.5-7B, and LLaVA-v1.6-Mistral-7B. To test across different scales and architectures, we applied VSFA to two additional models: Gemma 3 IT (4B) Team (2025) and Llama 3.2 Vision (11B) Team (2024). Gemma 3 IT uses a 4B language backbone with a SigLIP-based vision encoder. Llama 3.2 Vision uses an 11B backbone with a different vision-language integration architecture. Neither shares a vision encoder, projection layer, or LLM backbone with the Qwen or LLaVA families. We ran the standard VSFA pipeline on each model with only 60 training images, fine-tuned with LoRA under identical settings, and evaluated on FigStep with GPT-4o as the judge. Table 12 shows the results.
Table 12: VSFA on additional model families. Each model is trained with only 60 images. Lower ASR means safer.
Model Size Baseline ASR VSFA ASR Diff
Gemma 3 IT 4B 37.0% 7.2% -29.8%
Llama 3.2 Vision 11B 49.8% 17.6% -32.2%
Gemma 3 IT drops from 37.0% to 7.2% ASR, and Llama 3.2 Vision drops from 49.8% to 17.6%. Both are large reductions with only 60 training images. All six models now tested use different vision encoders and LLM backbones, covering sizes from 4B to 11B. VSFA produces consistent ASR reductions across all of them. This suggests that the mechanism is not tied to any particular architecture or model size. It is consistent with our SAE finding (Appendix C) that VSFA activates general safety-oriented persona features that appear to exist across different model families. The data efficiency is also notable. Our main experiments use 700 images, but here we used only 60 and still achieved large ASR reductions. This confirms that VSFA does not require large-scale training data to be effective.